OpenStudy (rational):

show that \(x^n-y^n\) is divisible by \(x-y\) for all \(n\in \mathbb{Z^+}\) using contradiction \(x,y\) are distinct integers

OpenStudy (anonymous):

IDK how haha

OpenStudy (rational):

since this is a proof by contradiction i think we start by assuming the opposite is true : suppose there exists some integer \(k\gt 0\) for which the given statement does not hold \[\dfrac{x^k-y^k}{x-y} \not \in \mathbb{Z}\]

OpenStudy (anonymous):

if x =y than x - y = 0 and we have error ))

OpenStudy (rational):

thats a good catch, i think the question requires some modification

OpenStudy (rational):

ive updated... see if it looks fine nw :)

OpenStudy (rational):

here im trying to mimic the proof of sqrt(2) irrationality, specifically this infinite descent argument http://en.wikipedia.org/wiki/Proof_by_infinite_descent#Irrationality_of_.E2.88.9A2

OpenStudy (rational):

im still working on proof using contradiction by infinite descent at the moment but i like to see other ways using contradiction too

OpenStudy (michele_laino):

let's suppose that exists an integer n_0 such that: \[\large{x^{{n_0}}} - {y^{{n_0}}} = r\left( {x - y} \right) + q\] where \[\large \begin{gathered} r = r\left( {x,y} \right) \hfill \\ q = q\left( {x,y} \right) \hfill \\ \end{gathered} \] now we can write: \[\large \begin{gathered} {x^{2{n_0}}} - {y^{2{n_0}}} = \left( {{x^{{n_0}}} - {y^{{n_0}}}} \right)\left( {{x^{{n_0}}} + {y^{{n_0}}}} \right) = \hfill \\ = \left( {r\left( {x - y} \right) + q} \right)\left( {{x^{{n_0}}} + {y^{{n_0}}}} \right) = \hfill \\ = r\left( {x - y} \right)\left( {{x^{{n_0}}} + {y^{{n_0}}}} \right) + q\left( {{x^{{n_0}}} + {y^{{n_0}}}} \right) \hfill \\ \end{gathered} \] so also \[\large{x^{2{n_0}}} - {y^{2{n_0}}}\] is not divisible by (x-y) and also: \[\large{x^{4{n_0}}} - {y^{4{n_0}}}\] is not divisible by (x-y) and so on...

OpenStudy (rational):

That looks neat! but that only works for even powers right ?

OpenStudy (michele_laino):

yes!

OpenStudy (rational):

your proof actually shows that if the given statemetn doesn't work for some \(n\), then it also doesn't work for any even multiple of \(n\) pretty amazing!!!

OpenStudy (michele_laino):

thanks!

OpenStudy (dan815):

http://prntscr.com/6q7fbq

OpenStudy (dan815):

it says how come we can say that x^n0 + y^n0 is not divisible by (x-y) for all n

OpenStudy (rational):

thats a good catch, i think we can fix it easily by using long division

OpenStudy (rational):

Here is the complete proof using michele's idea+long division + infinite descent :

OpenStudy (rational):

suppose there exists some integer \(k\gt 0\) for which the given statement doesn't hold. then : \[\dfrac{x^k-y^k}{x-y} \not \in \mathbb{Z} \] By long division we get \[\dfrac{x^k-y^k}{x-y} = x^{k-1} + \left(\color{blue}{\dfrac{x^{k-1}-y^{k-1}}{x-y}} \right)y \] Since \(x^{k-1}\) and \(y\) are integers, the blue part above must not be a integer for the left hand side to be a noninteger : \[\color{blue}{\dfrac{x^{k-1}-y^{k-1}}{x-y}} \not \in \mathbb{Z}\] Keep doing this till we reach \[\color{blue}{\dfrac{x^{1}-y^{1}}{x-y}} \not \in \mathbb{Z}\] Contradiction. \(\blacksquare\)

OpenStudy (dan815):

oo :)

OpenStudy (michele_laino):

I think that we can say that x^n_0+y^n0 is not divisible by x-y, using the mathematics induction principle @dan815

OpenStudy (rational):

i think we can put every induction proof in this infinite descent form

OpenStudy (rational):

i mean we can translate every induction proof into a contradiction proof using infinite descent

OpenStudy (kainui):

Wouldn't just directly factoring this work as well?

OpenStudy (rational):

thats how we do it normally, this thread is dedicated to infinite descent technique though ;p